The permiter of a rectangle is twice the sum of it's length and its width. The perimiter is 34 meters and its length is is 2 meters more than twice its length.
What is the length?

Answer by stanbon(75887)

The perimeter of a rectangle is twice the sum of it's length and its width. The perimeter is 34 meters and its length is is 2 meters more than twice its length.
What is the length?
---------------------------
P = 2(L + W)
---------
34 = 2(2W+2 + W)
-------
17 = 3W + 2
-------
3W = 15
-------
Width = 5 meters
------------------------
Length = 2W+2 = 12 meters
===========================
Cheers,
Stan H.
===========================